OpenStudy (anonymous):

Use this formula to solve this problem: Power (P) = Work (W) ÷ time (t) Joachim's body supplies a force of 450 N to run up a 20 meter hill in 15 seconds. How much power is involved in Joachim's run up the hill? A. 30 Watts B. 22.5 Watts C. 300 Watts D. 600 Watts

OpenStudy (abb0t):

\(\sf \color{red}{HINT}\): W = F\(\times\)d F = force [N] d = distance [m]

OpenStudy (abb0t):

That means you have: \(\sf \large \frac{F \times d}{t}\)